OpenStudy (anonymous):

Let f(x) = x2 - 3x - 7. Find f(-3).

OpenStudy (anonymous):

If f(x)=x^2 - 3x - 7 Then x=-3. -3^2 - 3(-3) - 7 9 --9 - 7 9+9-7 18-7 f(-3)=11

OpenStudy (anonymous):

You can also do it using synthetic substitution if you would like to see that as well?